A ball is left to drop, from rest, vertically from the top of a building from a position 48 m above ground.
i. What is the distance travelled by the ball after it has been falling for 3 seconds?
ii. What is the downwards velocity of the ball as it strikes the ground?
1
Expert's answer
2019-07-18T07:52:00-0400

The ball is moving starting from rest with constant acceleration "g = 9.8\, \text{m/s}^2". The distance "s" and velocity "v" after time "t" are given, respectively, by the relations

"s = \frac12 g t^2 \, , \qquad v = g t \, ."

Expressing the time "t = \sqrt{2 s/g}" from the first relation, and substituting it into the second one, we obtain the velocity as a function of the distance:

"v = \sqrt{2 g s} \, ."

Substituting "t = 3\, \text{s}" into the first relation, and "s = h = 48\, \text{m}" into the last one, we obtain the answers to the problem:


i. "s = 44.1\, \text{m}"

ii. "v = 30.7\, \text{m/s}"